# Retrieval-Enhanced Machine Learning: Synthesis and Opportunities

### v1

@@ Abstract.
journal: CSUR
@@ 1.1. Background
The convergence of these two domains has given rise to a new trend in research, where models are equipped with retrieval capabilities to access external knowledge during both training and inference stages (Mialon et al., 2023; Zamani et al., 2022). This integration of retrieval mechanisms into the prediction pipeline started to gain significant traction, as it allows models to ground their predictions in external knowledge without necessitating an increase in model capacity. Methods presented by Hashemi et al. (2020) and Lewis et al. (2020b) are among the earliest work in this space; the former focuses on retrieval-augmented representation learning by extending the transformer network, while the latter studies the paradigm of retrieval-augmented generation (RAG) for knowledge-intensive language tasks. That said, using retrieval results to improve a machine learning systems is not new. Pseudo-relevance feedback methods—methods for representing search queries using the top retrieved documents—are perhaps the first set of methods in this category (Attar and Fraenkel, 1977; Croft and Harper, 1979). The ICL ability inherent in the LLMs has played a pivotal role in facilitating the dissemination and adoption of these retrieval-augmented approaches. By integrating retrieved documents into the prompt of the LLMs, researchers have been able to harness the external knowledge sources without fundamentally altering the underlying model architecture.
1.2. Motivation
@@
Since improving model performance by increasing the number of parameters is not sustainable, one motivation of retrieval-based models stems from the finding that, while large models tend to memorize training data (Carlini et al., 2021), incorporating retrieval-based methods can effectively transfer the burden of memorization to external storage systems (Borgeaud et al., 2022). We advocate for enhancing machine learning (ML) models in general (i.e., beyond generation) with the ability to employ stored information via information retrieval techniques. IR has already shown its merits in aiding human interaction with vast text databases. We posit that IR’s utility can be broadened to enable machine access to not only extensive text databases but also to knowledge represented in more abstract forms. By integrating ML architectures with direct access to IR systems, we aim to separate the processes of reasoning and memory. Zamani et al. (2022) dubbed this approach, retrieval-enhanced machine learning (REML), as a broader concept that extends ML. Extending their work, we further survey the recent advances of REML in the field of ML, including NLP, with consistent mathematical notation. By doing so, we aim to equip researchers with a comprehensive and structured overview of the REML methodologies, enabling them to swiftly embark on research within this domain.
1.3. Applications of REML
@@
The landscape of REML paradigm encompasses a diverse array of sub-domains, each with its unique set of challenges and applications. This includes seminal work in language modeling (Guu et al., 2020; Lewis et al., 2020b; Borgeaud et al., 2022; Izacard and Grave, 2021b; Zhong et al., 2022; Izacard et al., 2024; Ram et al., 2023; Wang et al., 2023b; Shi et al., 2024; Li et al., 2022; Khandelwal et al., 2020; Lyu et al., 2023), machine translation (Khandelwal et al., 2021), question answering (Yu et al., 2022a; Chen et al., 2017a; Lee et al., 2019; Nakano et al., 2022; Lazaridou et al., 2022; Wu et al., 2022d; Chen et al., 2023c, d; Zhang et al., 2024), fact verification (Lewis et al., 2020b; Petroni et al., 2023; Chen et al., 2023b), open domain (Shuster et al., 2022b, a; Komeili et al., 2022; Thoppilan et al., 2022) and task-oriented (Thulke et al., 2021; Cai et al., 2023; Eric et al., 2017; Raghu et al., 2021; Nekvinda and Dušek, 2022) dialog systems, slot filling (Glass et al., 2021), state tracking (King and Flanigan, 2023), multimodal dialog (Fan et al., 2021), reinforcement learning (Fernández and Veloso, 2006; Goyal et al., 2022; Humphreys et al., 2022), computer vision (Chen et al., 2023a; Yasunaga et al., 2023; Ramos et al., 2023; Shrestha et al., 2024), commonsense reasoning (Yu et al., 2022c), evidence attribution (Gao et al., 2023b; Aksitov et al., 2023; Menick et al., 2022; Huo et al., 2023; Gao et al., 2023a), knowledge-graph augmentation (Sen et al., 2023; Kang et al., 2023; Baek et al., 2023; Ju et al., 2022; Hu et al., 2023; Zhang et al., 2022b; Yu et al., 2022a), ranking (Hui et al., 2022a), personalization (Salemi et al., 2024b, a), mathematical problem-solving (Yang et al., 2024b), code generation (Zhang et al., 2023; Zhou et al., 2023; Wang et al., 2024a), representation learning for audio and speech (Sanabria et al., 2023; Lin et al., 2024), time series forecasting (Jing et al., 2022; Yang et al., 2022), and protein structure prediction (Cramer, 2021). The industry and open source communities have swiftly embraced the adoption of retrieval-based models, recognizing their potential for accelerated adaptation and performance enhancement. Frameworks such as LangChain, [^1] https://www.langchain.com LlamaIndex, [^2] https://www.llamaindex.ai and DSPy (Khattab et al., 2024) have emerged, streamlining the process of implementing the retrieval-based models. This broad spectrum of domains (not an exhaustive list) underscores the versatility and impact of the REML paradigm across diverse applications.
1.4. Main Contributions of This Work
@@
Although many current applications are centered around natural language processing, we believe that ML models that leverage retrieval components are not confined to language models alone, but can be extended to any ML models. To address this broader applicability, we formalize the framework as Retrieval-Enhanced Machine Learning (REML) and synthesize existing studies with consistent mathematical notations which is lacking in the current literature. Also, despite the advancements in REML models, there remains a significant underutilization of the rich and extensive body of work from information retrieval research which can offer numerous methodologies and insights that can substantially benefit REML models. This work aims to bridge this gap by integrating IR research into the design of REML models. Ultimately, we hope this work will enable researchers across various fields leveraging ML to easily understand the framework of REML and its extensibility.
@@ 2. Retrieval-Enhanced Machine Learning
[…]
@@ 3. Querying
Querying refers to how requests are represented and constructed. how do we define $\mathcal{Q}$? how does $f_{\theta}$ construct $q\in\mathcal{Q}$ from $x$?
@@ 3.1. Deciding Where to Query
Before sending a query to an information access system, a REML system can decide where its query should be sent to. Bearing in mind that each $g_{\omega_{i}}$ is associated with $\textrm{C}_{i}$ (multiplicative nature of information access process depicted in Figure 2), the Query Generation module first decides which tuple(s) of $g_{\omega_{i}}$ and $\textrm{C}_{i}$ should be selected depending on the context which aligns with the mixture-of-expert-like interpretation described by Pan et al. (2023)  [^3] In KiC (Pan et al., 2023), a router selects an expert predictive model (not a retriever) with specific knowledge source. However, we think that this helps the understanding of the first step of the Query Generation module.. This query decision problem can be understood by the following sub-problems. 1) Corpus Selection: deciding which corpora are needed to be searched over (can be null when no retrieval is needed), and 2) Retriever Selection: deciding which retrieval model should be used to search over the chosen corpora.
@@ 3.1.1. Corpus Selection
Selecting $\textrm{C}_{i}$ is not only regarded with deciding what kind of external information should be provided to the $f_{\theta}$ (Pan et al., 2023) but also encompasses the question of when to query (no corpus selection when no retrieval is beneficial to $f_{\theta}$). This can be a critical question as retrieval augmentation can hurt the performance for certain input types (Maekawa et al., 2024; Mallen et al., 2023; Asai et al., 2024). It also can save computational resources by reducing the number of searching (Labruna et al., 2024). There can be several criteria on whether external information can be beneficial to the predictive model. It can be based on term popularity (Mallen et al., 2023), input complexity (Jeong et al., 2024), or a trained model (Asai et al., 2024).
@@ 3.2.2. Expansion
In certain scenarios, the input alone may lack essential information required by the search system to yield desired results. In such situations, augmenting the input with additional pertinent data can be done. This process of expansion broadens the context and enhances the search system’s capability to retrieve relevant and meaningful results, thereby improving overall system performance. Typically, expansion is achieved by concatenating the input with previously retrieved results (Zhu et al., 2021; Xiong et al., 2021) or generated text that is conditioned on the input (Wang et al., 2023d; Mao et al., 2021; Liu et al., 2022; Chuang et al., 2023).
@@ 3.4. Unified Equation for Query Generation
Any of these function can be replaced with the identity function (Izacard and Grave, 2021b; Karpukhin et al., 2020; Asai et al., 2022; Yamada et al., 2021; Lewis et al., 2020b, b; Thorne et al., 2018; Guu et al., 2020) disregarding modality (Salemi et al., 2023b; Qu et al., 2021). Furthermore, it is conceivable to apply these functions multiple times and in various order. Particularly in the realm of multi-hop question answering and fact verification, prior research extensively employs multiple transformations and decompositions to fulfill the task requirements (Qi et al., 2019; Trivedi et al., 2023; Yadav et al., 2020; Das et al., 2019; Jiang et al., 2023). Given the intricacies of these tasks, leveraging a combination of various transforms and decompositions becomes essential.
@@ 4. Searching
Searching refers to how requests and items are combined to construct the retrieval results. how does $g_{\omega}$ generate $r\in\mathcal{R}$ from $q\in\mathcal{Q}$
@@ 4.2. Retrieval Models with Dense Representations
Queries and documents can be represented using low-dimensional (compared to the vocabulary size) dense vectors. Such dense vectors are often obtained using pre-trained language models, such as BERT (Devlin et al., 2019), that are fine-tuned for retrieval tasks (Karpukhin et al., 2020). Dense retrieval models are commonly based on bi-encoder architectures – one encoder for representing the query and another one for representing the document. These encoders can share parameters. Some dense retrieval methods, such as DPR (Karpukhin et al., 2020), represent each query or document by a single dense vector. While others, such as ColBERT (Khattab and Zaharia, 2020), use one vector per token, resulting in multiple vectors for each query and document. Approximate nearest neighbor (ANN) algorithms, such as HNSW (Malkov and Yashunin, 2020), are used for efficient retrieval when dealing with dense representations. Dense retrieval approaches are also commonly used when dealing with multi-media and multi-modal data (Qu et al., 2021; Salemi et al., 2023a).
@@ 4.5.1. Content-based addressing
At $t$’th iteration, with slight abuse of notation (simplifying $\textrm{C}_{i}^{t}$ to $\textrm{C}_{t}$), given a query $q_{t}$, a collection $\textrm{C}_{t}$, content-based addressing can be defined as:
@@ 4.5.2. Location-based addressing
where combine is a function that generates an address based on the location-based and content-based addresses. Although most of the previous work uses either pure content-based addressing (Santoro et al., 2016; Guu et al., 2020; Khandelwal et al., 2020; Rae et al., 2016) or pure location-based addressing (Weston et al., 2015; Shinn et al., 2023), some work employs both content- and location-based addressing (Graves et al., 2014, 2016).
@@ 4.5.3. Unified Equation for Searching
With the final address vector $w_{t}$, the retrieval results $r_{t}$ from $g_{\omega}$ can be defined as
@@ 5. Presentation & Consumption
Presentation refers to how retrieval results are represented and consumption refers to how they are incorporated into the predictive model. how do we define $\mathcal{R}$? how does $f_{\theta}$ incorporate/use $r\in\mathcal{R}$?
@@ 5.1.1. Transforming the data
Due to possible context limits of predictive model, it is desirable to condense document data so that more documents can fit into the context. This can be achieved through automatic summarization, converting the original data into a shortened form through an extractive or abstractive process (Gao et al., 2023b; Li et al., 2023). Furthermore, data can be summarized in the context of the input, providing clarity and explaining why the document is relevant.
@@ 5.1.4. Unified Equation for Presentation
where any of these functions can be replaced with simple forms such as an identity function (Lewis et al., 2020a; Izacard and Grave, 2021b, a). Additionally, we can imagine these functions being applied multiple times in any order. We include this ordering as the one that seems most natural when taking into account context length limits. $transform_{p}$ is on individual items, and $compose$ is similar to $transform_{p}$ but on groups of items. Not shown is loading of retrieved items. Unlike traditional ML systems, where data is typically used only for training, REML systems have unique requirements associated with data loading. Since the amount of external data required for an input is dynamic and considerable (Borgeaud et al., 2022), efficient load and manage of data is essential (Douze et al., 2024; Guo et al., 2020b). We assume loading is handled implicitly by the retrieval module.
Transform (§5.1.1) ALCE (Gao et al., 2023b) Explores both summarization and extractive snippets to compress retrieved items. Teach LLMs to Personalize (Li et al., 2023) Context independent and dependent summarization to emphasize key retrieved aspects. QADecontext (Newman et al., 2023) $\dagger$ An example application where decontextualization is done as the downstream task when presenting passages from scientific documents. Compose (§5.1.2) Fixed Chunking (Wu et al., 2021) Recursively summarizes adjacent chunks in books. RAPTOR (Sarthi et al., 2024) First clusters then summarizes related chunks of text. Truncate (§5.1.3) FiD-Light (Hofstätter et al., 2023) Extracts subset lists of vectors in an FiD-like model to speed up attention-related performance bottleneck during decoding. Choppy (Bahri et al., 2020) $\dagger$ A supervised approach to ranked list truncation.

### v2

@@ Abstract.
Information Retrieval, Machine Learning
ccs: Information systems Information retrievalccs: Computing methodologies Machine learning
@@ 1. Introduction
The convergence of these two domains has given rise to a new trend in research, where models are equipped with retrieval capabilities to access external knowledge during both training and inference stages (Mialon et al., 2023; Zamani et al., 2022). This integration of retrieval mechanisms into the prediction pipeline started to gain significant traction, as it allows models to ground their predictions in external knowledge without necessitating an increase in model capacity. Methods presented by Hashemi et al. (2020) and Lewis et al. (2020) are among the earliest work in this space; the former focuses on retrieval-augmented representation learning by extending the transformer network, while the latter studies the paradigm of retrieval-augmented generation (RAG) for knowledge-intensive language tasks. That said, using retrieval results to improve a machine learning systems is not new. Pseudo-relevance feedback methods—methods for representing search queries using the top retrieved documents—are perhaps the first set of methods in this category (Attar and Fraenkel, 1977; Croft and Harper, 1979). The ICL ability inherent in the LLMs has played a pivotal role in facilitating the dissemination and adoption of these retrieval-augmented approaches. By integrating retrieved documents into the prompt of the LLMs, researchers have been able to harness the external knowledge sources without fundamentally altering the underlying model architecture.
Motivation
@@
Since improving model performance by increasing the number of parameters is not sustainable, one motivation of retrieval-based models stems from the finding that, while large models tend to memorize training data (Carlini et al., 2021), incorporating retrieval-based methods can effectively transfer the burden of memorization to external storage systems (Borgeaud et al., 2022; Shao et al., 2024). We advocate for enhancing machine learning (ML) models in general (i.e., beyond generation) with the ability to employ stored information via information retrieval techniques. IR has already shown its merits in aiding human interaction with vast text databases. We posit that IR’s utility can be broadened to enable machine access to not only extensive text databases but also to knowledge represented in more abstract forms. By integrating ML architectures with direct access to IR systems, we aim to separate the processes of reasoning and memory. Zamani et al. (2022) dubbed this approach, retrieval-enhanced machine learning (REML), as a broader concept that extends ML. Extending their work, we further survey the recent advances of REML in the field of ML, including NLP, with consistent mathematical notation. By doing so, we aim to equip researchers with a comprehensive and structured overview of the REML methodologies, enabling them to swiftly embark on research within this domain.
Applications of REML
@@
The landscape of REML paradigm encompasses a diverse array of sub-domains, each with its unique set of challenges and applications. This includes seminal work in language modeling (Guu et al., 2020; Lewis et al., 2020; Borgeaud et al., 2022; Izacard and Grave, 2021b; Zhong et al., 2022; Izacard et al., 2024; Ram et al., 2023; Wang et al., 2023c; Shi et al., 2024; Li et al., 2022; Khandelwal et al., 2020), machine translation (Khandelwal et al., 2021), question answering (Yu et al., 2022a; Chen et al., 2017a; Lee et al., 2019; Nakano et al., 2022; Lazaridou et al., 2022; Wu et al., 2022d; Chen et al., 2023b; Zhang et al., 2024a; Kim et al., 2023), fact verification (Lewis et al., 2020; Petroni et al., 2023; Chen et al., 2024), open domain (Shuster et al., 2022; Komeili et al., 2022; Thoppilan et al., 2022) and task-oriented (Thulke et al., 2021; Eric et al., 2017; Raghu et al., 2021; Nekvinda and Dušek, 2022) dialog systems, slot filling (Glass et al., 2021), state tracking (King and Flanigan, 2023), reinforcement learning (Fernández and Veloso, 2006; Goyal et al., 2022; Humphreys et al., 2022; Chowdhury et al., 2024), computer vision (Chen et al., 2023a; Yasunaga et al., 2023; Ramos et al., 2023; Shrestha et al., 2024), multimodal ML (Fan et al., 2021; Iscen et al., 2024), commonsense reasoning (Yu et al., 2022c), evidence attribution (Gao et al., 2023b; Menick et al., 2022; Huo et al., 2023; Gao et al., 2023a), knowledge-graph augmentation (Baek et al., 2023; Ju et al., 2022; Hu et al., 2023; Zhang et al., 2022b; Yu et al., 2022a; Wei et al., 2024), personalization (Salemi et al., 2024b, a), mathematical problem-solving (Yang et al., 2024b), code generation (Zhang et al., 2023; Zhou et al., 2023; Wang et al., 2024a), representation learning for audio and speech (Sanabria et al., 2023; Lin et al., 2024; Zhou et al., 2024), time series prediction (Jing et al., 2022; Yang et al., 2022; Xu et al., 2024a), robot navigation and embodiment (Xie et al., 2024; Anwar et al., 2024), chip design (Chowdhury et al., 2024), medical reasoning (Jeong et al., 2024b), drug discovery (Liu et al., 2024), and protein structure prediction (Jumper et al., 2021; Cramer, 2021). The industry and open source communities have swiftly embraced the adoption of retrieval-based models, recognizing their potential for accelerated adaptation and performance enhancement. Frameworks such as LangChain, [^1] https://www.langchain.com LlamaIndex, [^2] https://www.llamaindex.ai and DSPy (Khattab et al., 2024) have emerged, streamlining the process of implementing the retrieval-based models. This broad spectrum of domains (not an exhaustive list) underscores the versatility and impact of the REML paradigm across diverse applications.
Main Contributions of This Work
@@
Although many current applications are centered around natural language processing, we believe that ML models that leverage retrieval components are not confined to language models alone, but can be extended to any ML models. To address this broader applicability, we formalize the framework as Retrieval-Enhanced Machine Learning (REML) and synthesize existing studies with consistent mathematical notations which is lacking in the current literature. Throughout this paper, we consistently use the notations described in Table 1. Also, despite the advancements in REML models, there remains a significant underutilization of the rich and extensive body of work from information retrieval research which can offer numerous methodologies and insights that can substantially benefit REML models. This work aims to bridge this gap by integrating IR research into the design of REML models. Ultimately, we hope this work will enable researchers across various fields leveraging ML to easily understand the framework of REML and its extensibility.
@@ 2. Retrieval-Enhanced Machine Learning
Notation Description Notation Description $x$ input instance $d$ retrieval item $\mathcal{X}$ input space $\mathcal{D}$ retrieval space $y$ output target $\mathcal{D}_{\text{key}}$ retrieval key space $\mathcal{Y}$ output space C stored retrievable items $q$ query L labeled data (i.e., $\textrm{L}\subset\mathcal{X}\times\mathcal{Y}$) $\mathcal{Q}$ query space U unlabeled data (i.e., $\textrm{U}\subset\mathcal{X}$) $r$ retrieval results $\mathcal{R}$ retrieval result space $\mathcal{L}$ The downstream loss function $s$ model feedback $f_{\theta}$ a predictive machine learning model parameterized by $\theta$ $\mathcal{S}$ model feedback space $g_{\omega}$ a retrieval model parameterized by $\omega$ $\mu$ evaluation metric $\mathcal{M}$ a tuple of predictive and retrieval model
@@ 3.1. Deciding Where to Query
Before sending a query to an information access system, a REML system can decide where its query should be sent to. Bearing in mind that each $g_{\omega_{i}}$ is associated with $\textrm{C}_{i}$ (multiplicative nature of information access process depicted in Figure 2), the Query Generation module first decides which tuple(s) of $g_{\omega_{i}}$ and $\textrm{C}_{i}$ should be selected depending on the context which aligns with the mixture-of-expert-like interpretation described by Pan et al. (2023)  [^3] In KiC (Pan et al., 2023), a router selects an expert predictive model (not a retriever) with specific knowledge source. However, this helps the understanding of the first step of the Query Generation module.. This query decision problem can be understood by the following sub-problems. 1) Corpus Selection: deciding which corpora are needed to be searched over (can be null when no retrieval is needed), and 2) Retriever Selection: deciding which retrieval model should be used to search over the chosen corpora.
@@ 3.1.1. Corpus Selection
Selecting $\textrm{C}_{i}$ is not only regarded with deciding what kind of external information should be provided to the $f_{\theta}$ (Pan et al., 2023) but also encompasses the question of when to query (no corpus selection when no retrieval is beneficial to $f_{\theta}$). This can be a critical question as retrieval augmentation can hurt the performance for certain input types (Maekawa et al., 2024; Mallen et al., 2023; Asai et al., 2024). It also can save computational resources by reducing the number of searching (Labruna et al., 2024). There can be several criteria on whether external information can be beneficial to the predictive model. It can be based on term popularity (Mallen et al., 2023), input complexity (Jeong et al., 2024a), or a trained model (Asai et al., 2024).
@@ 3.2.2. Expansion
In certain scenarios, the input alone may lack essential information required by the search system to yield desired results. In such situations, augmenting the input with additional pertinent data can be done. This process of expansion broadens the context and enhances the search system’s capability to retrieve relevant and meaningful results, thereby improving overall system performance. Typically, expansion is achieved by concatenating the input with previously retrieved results (Zhu et al., 2021; Xiong et al., 2021) or generated text that is conditioned on the input (Wang et al., 2023e; Mao et al., 2021; Liu et al., 2022; Chuang et al., 2023).
@@ 3.4. Unified Equation for Query Generation
Any of these function can be replaced with the identity function (Izacard and Grave, 2021b; Karpukhin et al., 2020; Asai et al., 2022; Yamada et al., 2021; Lewis et al., 2020, 2020; Thorne et al., 2018; Guu et al., 2020) disregarding modality (Salemi et al., 2023b; Qu et al., 2021). Furthermore, it is conceivable to apply these functions multiple times and in various order. Particularly in the realm of multi-hop question answering and fact verification, prior research extensively employs multiple transformations and decompositions to fulfill the task requirements (Qi et al., 2019; Trivedi et al., 2023; Yadav et al., 2020; Das et al., 2019; Jiang et al., 2023). Given the intricacies of these tasks, leveraging a combination of various transforms and decompositions becomes essential.
@@ 4.2. Retrieval Models with Dense Representations
Queries and documents can be represented using low-dimensional (compared to the vocabulary size) dense vectors. Such dense vectors are often obtained using pre-trained language models, such as BERT (Devlin et al., 2019), that are fine-tuned for retrieval tasks (Karpukhin et al., 2020). Dense retrieval models are commonly based on bi-encoder architectures – one encoder for representing the query and another for the document. These encoders could share parameters. Some dense retrieval methods, such as DPR (Karpukhin et al., 2020), represent each query or document by a single dense vector. While others, such as ColBERT (Khattab and Zaharia, 2020), use one vector per token, resulting in multiple vectors for each query and document. Approximate nearest neighbor (ANN) algorithms, such as HNSW (Malkov and Yashunin, 2020), are used for efficient retrieval when dealing with dense representations. Dense retrieval approaches are also commonly used when dealing with multi-media and multi-modal data (Qu et al., 2021; Salemi et al., 2023a).
@@ 4.5.1. Content-based addressing
At $t$’th iteration, with a slight simplification of the notation (simplifying $\textrm{C}_{i}^{t}$ to $\textrm{C}_{t}$), given a query $q_{t}$ and a collection $\textrm{C}_{t}$, content-based addressing can be defined as:
@@ 4.5.2. Location-based addressing
where combine is a function that generates an address based on the location-based and content-based addresses. For example, most previous work relies on either purely content-based addressing (Santoro et al., 2016; Guu et al., 2020; Khandelwal et al., 2020; Rae et al., 2016) or purely location-based addressing (Weston et al., 2015; Shinn et al., 2023). However, it is also possible to combine both content-based and location-based addressing (Graves et al., 2014, 2016).
@@ 4.5.3. Unified Equation for Searching
With the final address vector $w_{t}$, the retrieval results $r_{t}$ from $g_{\omega}$ is defined as:
@@ 5.1.1. Transforming the data
Due to possible context limits of predictive model, it is desirable to condense document data so that more documents can fit into the context. This can be achieved through automatic summarization, converting the original data into a shortened form through an extractive or abstractive process (Gao et al., 2023b; Li et al., 2023; Kim et al., 2024; Xu et al., 2024b). Furthermore, data can be summarized in the context of the input, providing clarity and explaining why the document is relevant.
@@ 5.1.4. Unified Equation for Presentation
where any of these functions can be replaced with simple forms such as an identity function (Lewis et al., 2020; Izacard and Grave, 2021b, a). Additionally, we can imagine these functions being applied multiple times in any order. We include this ordering as the one that seems most natural when taking into account context length limits. $transform_{p}$ is on individual items, and $compose$ is similar to $transform_{p}$ but on groups of items. Not shown is loading of retrieved items. Unlike traditional ML systems, where data is typically used only for training, REML systems have unique requirements associated with data loading. Since the amount of external data required for an input is dynamic and considerable (Borgeaud et al., 2022), efficient load and manage of data is essential (Douze et al., 2024; Guo et al., 2020b). We assume loading is handled implicitly by the retrieval module.
Transform (§5.1.1) ALCE (Gao et al., 2023b), SuRe (Kim et al., 2024) & RECOMP (Xu et al., 2024b) Explore summarization of retrieved items for compression. Teach LLMs to Personalize (Li et al., 2023) Context independent and dependent summarization to emphasize key retrieved aspects. QADecontext (Newman et al., 2023) $\dagger$ Decontextualization as a downstream task when presenting passages from scientific documents. Compose (§5.1.2) Fixed Chunking (Wu et al., 2021) Recursively summarizes adjacent chunks in books. RAPTOR (Sarthi et al., 2024) First clusters then summarizes related chunks of text. Truncate (§5.1.3) FiD-Light (Hofstätter et al., 2023) Extracts vector subsets to speed up attention bottlenecks in FiD-like model decoding. Choppy (Bahri et al., 2020) $\dagger$ A supervised approach to ranked list truncation.